# suppSupplementary Material References

### v1

@@ 4.1 Signed Stochastic Block Model Approximation
For any model specification, we recover the Signed Stochastic Block Model (SSBM), as shown in Example 1, by setting all parameters in model (4) corresponding to dyad-dependent statistics (e.g., transitivity) to zero. Note that the sufficient statistics in (3) are already dyad-independent by definition. Therefore, the approximation of the probability of our model by the probability of a SSBM only accrues errors for within-block networks. As $N$ and $K$ grow, our joint probability will then be increasingly dominated by terms relating to between-block probabilities, provided single blocks are not too large. This offers a heuristic justification for why the SSBM provides a good approximation in large populations.
@@
All sufficient statistics considered in this paper satisfy Assumptions 4.1 and 4.1. Following Theorem 2 from Babkin et al. (2020), the full conditional probability $\mathbb{P}_{\bm{\theta}}(\mathbf{Y}=\mathbf{y}\mid\bm{Z}=\bm{z})$ can be then approximated with $\bm{\theta}_{SBM}$ as follows
@@ 4.2 Variational Approximation based on MM Updates
Supported by this result, we substitute $\mathbb{P}_{\bm{\theta}}(\mathbf{Y}=\mathbf{y}\mid\bm{Z}=\bm{z})$ in (8) by the probability of a SSBM, which simplifies the optimization in Steps 1 and 2 outlined above. We assume a mean-field variational family for the class of discrete distributions characterized, implying that the joint distribution of $\bm{Z}$ factorizes across nodes:
@@
4.3 Parameter Step with known Blocks
@@
$\begin{array}[]{llllllllll}\hat{\bm{z}}=\left(\operatorname*{arg\,max}\limits_{k\,\in\,\{1,...,K\}}\alpha_{i,k}\right)_{i=1}^{N}.\end{array}$
@@
The term $\bm{a}\otimes\bm{b}$ denotes the Kronecker product of vectors $\bm{a}\in\mathbb{R}^{m}$ and $\bm{b}\in\mathbb{R}^{n}$. In the following, we focus on estimating the within-block model parameter $\bm{\beta}_{w,\operatorname{vec}}$. The method applies analogously to $\bm{\beta}_{b,\operatorname{vec}}$. While both MCMC-based maximum likelihood (Hunter and and Handcock, 2006) and maximum pseudo-likelihood methods (MPLE, Strauss and and Ikeda, 1990) can be used to estimate the parameters in the second step, MPLE is generally better suited for large networks due to its computational efficiency and scalability. If dyadic independence holds, both approaches coincide.
@@
With $\Delta_{i,j,k}^{0\rightarrow+}(\mathbf{y})$ and $\Delta_{i,j,k}^{0\rightarrow-}(\mathbf{y})$ being the change in the reparametrized sufficient statistics given in (11) after changing the edge value $y_{i,j}$ from ``$0"$ to ``+"``"``+"` ` + " or ``"``"``-"` ` - ", respectively, we can express the relative log odds of observing $\mathbf{Y}_{i,j}$ to be ``+"``"``+"` ` + " abs ``"``"``-"` ` - " instead of ``$0"$, conditional on $z_{i}=z_{j}=k$:
$\begin{array}[]{llllllllll}\log\left(\dfrac{\mathbb{P}_{\bm{\beta}_{w,\operatorname{vec}}}(Y_{i,j}=``+"\mid\mathbf{Y}_{(-ij)}=\mathbf{y}_{(-ij)},z_{i}=z_{j}=k)}{\mathbb{P}_{\bm{\beta}_{w,\operatorname{vec}}}(Y_{i,j}=``0"\mid\mathbf{Y}_{(-ij)}=\mathbf{y}_{(-ij)},z_{i}=z_{j}=k)}\right)&=&\bm{\beta}_{w,\operatorname{vec}}^{\top}\Delta_{i,j,k}^{0\rightarrow+}(\mathbf{y})\vspace{0.25cm}\\ \log\left(\dfrac{\mathbb{P}_{\bm{\beta}_{w,\operatorname{vec}}}(Y_{i,j}=``-"\mid\mathbf{Y}_{(-ij)}=\mathbf{y}_{(-ij)},z_{i}=z_{j}=k)}{\mathbb{P}_{\bm{\beta}_{w,\operatorname{vec}}}(Y_{i,j}=``0"\mid\mathbf{Y}_{(-ij)}=\mathbf{y}_{(-ij)},\bm{z})}\right)&=&\bm{\beta}_{w,\operatorname{vec}}^{\top}\Delta_{i,j,k}^{0\rightarrow-}(\mathbf{y}).\end{array}$
@@
Newton Raphson algorithms are commonly used to maximize (12) based on (13) with the following gradient and negative Hessian:
@@ 4.4 Uncertainty Quantification
Given the previously sampled block memberships, $\bm{z}^{(1)},\ldots,\bm{z}^{(T)}$, the parameters of the model,$(\bm{\beta}_{w,\operatorname{vec}}^{(1)},\mbox{\boldmath$\Sigma$}^{(1)}),\ldots,(\bm{\beta}_{w,\operatorname{vec}}^{(T)},\mbox{\boldmath$\Sigma$}^{(T)})$, are re-estimate for each sampled partition to obtain valid uncertainty quantification. The estimated parameters $\bm{\beta}_{w,\operatorname{vec}}^{(t)}$ are pooled as follows
@@ 5 Simulation Study
This section evaluates the performance of the proposed estimation procedure in both steps. For the evaluation of the parameter recovery in Step 2, we use MPLE as they are more scalable than MCMC-based methods. To evaluate the block recovery we use the variational approach described in Section 4 and compare it to the spectral clustering described in Lei and Rinaldo (2015), for which we treat the network as a binary one. The block recovery will be assessed using the Yule’s $\phi$-coefficient:
@@ 6 Wikipedia Network
We demonstrate the proposed model in an application to a network of Wikipedia editors who contribute to and refine content across a range of topics. Wikipedia is a free, community-edited online encyclopedia that has been the subject of numerous prior studies (Brandes et al., 2009; Iba et al., 2010). A subset of these editors can be considered experts in specific fields, as they focus their contributions on a limited number of Wikipedia pages within their area of expertise. These editors typically exhibit limited awareness of edits or interactions occurring outside their domain. Such behavior gives rise to block structure and localized dependence within the network, where each block corresponds to a distinct area of knowledge.
@@ 6.1 Model Specification
Figure 3: Out-of-sample cross-validation results for the Full Triad model including all triadic terms ($\text{Edges}^{+/-}$, $\text{GWD}^{+/-}$, and all GWESP). The distribution of simulated statistics across 100 replications is compared against the observed statistics for each block.
@@
$\begin{array}[]{llllllllll}\bm{s}(\mathbf{y}_{k,k})&=&(&\text{Edges}^{+}(\mathbf{y}_{k,k}),\,\text{Edges}^{-}(\mathbf{y}_{k,k}),\,\text{GWD}^{+}(\mathbf{y}_{k,k},0.2),\,\text{GWD}^{-}(\mathbf{y}_{k,k},0.2),\\ &{}&{}&\text{GWESE}^{+}(\mathbf{y}_{k,k},0),\,\text{GWESF}^{+}(\mathbf{y}_{k,k},0),\,\text{GWESE}^{-}(\mathbf{y}_{k,k},0),\\ &{}&{}&\text{GWESF}^{+}(\mathbf{y}_{k,k},0))^{\top}.\end{array}$
The results of the out-of-sample cross-validation, shown in Figure 3 and Supplement C.1, suggest that the Full Triad model, including edges, GWD and all GWESP terms, is better equipped to describe the Wikipedia network.
@@ 6.2 Results
The edges terms are modeled as a function of block size $\log(N_{k})$. The terms $\text{Edges}^{+}$ and $\text{Edges}^{-}$ represent the intercepts, which alone are not directly interpretable, as they would only apply to hypothetical blocks of size 1. Instead, the combination of intercept and block size effect reflects the expected edge probability within a given block. Across all models, the coefficients for positive edges are lower than for negative edges, indicating that negative edges are more frequent within blocks. The consistently negative coefficients for the block size terms demonstrates that larger blocks tend to be sparser, with edge probability decreasing as block size increases. This reflects the fact that the number of possible edges grows quadratically, while the number of actual edges grows at a slower pace. Adding GWD terms reduces the edge coefficients, suggesting that some variation in edge probability is explained by degree structure. The GWD coefficients are negative for both edge types in all models, indicating a general tendency against high-degree nodes. The effect is stronger for positive edges, implying that these edges are more evenly distributed and less likely to form hubs. As expected from structural balance theory, the $\text{GWESE}^{+}$ terms have positive and statistically significant coefficients in both the Full and Partial Triad model. This means that users are more likely to restore the work of another user if they share a common enemy. While still positive and significant, the other balanced triad represented by $\text{GWESF}^{+}$ is closer to zero, similar to the unbalanced triad $\text{GWESE}^{-}$. However, the term $\text{GWESF}^{-}$ is significantly larger than $\text{GWESF}^{+}$, which is not consistent with structural balance theory. This discrepancy suggests that structural balance theory may not fully apply to this expert network, For instance, undoing a friend-of-a-friend’s contribution could be more common due to task-related disagreements rather than group-based antagonism. Overall, the more complex models capture important network features, such as triadic relationships, and this is reflected in their lower AIC values, indicating better model fit despite increased complexity.
Our Wikipedia network analysis of thousands of subject matter experts serves an example where global dependence assumption seems unrealistic. Through the use of local dependence, we found some adherence to the behavioral norms predicted by structural balance theory. Specifically, users who have an enemy in common are more likely to restore each other’s contributions. This finding is consistent with Lerner and Lomi (2019).
@@ 7 Conclusion
Our proposed model for signed networks combines the strengths of stochastic block models and exponential random graph models, while being scalable to thousands of nodes. There are several directions for extending the proposed model. First, we assumed that the number of blocks, $K$, was known a priori. In practice, this may not hold, and methods to infer $K$ from data would enhance model flexibility (Saldaña et al., 2017). Second, the assumption of nonoverlapping community structure could be relaxed by allowing nodes to belong to multiple blocks, as in mixed-membership models (see, e.g., Latouche et al., 2011). This would accommodate for more realistic block structures. Third, the observed degree heterogeneity in the Wikipedia network highlights the need for a degree-corrected stochastic block model for signed networks.
@@ B.1 Surrogate Function
$\begin{array}[]{llllllllll}Q({\mbox{\boldmath$\gamma$}}^{(t)},{\bm{\theta}}^{(t)};{\bm{\alpha}}^{(t)},{\bm{\alpha}})&=&\displaystyle\sum\limits_{i<j}\displaystyle\sum\limits_{k=1}^{K}\displaystyle\sum\limits_{l=1}^{K}\left(\alpha_{i,k}^{2}\dfrac{\alpha_{j,l}^{(t)}}{2\alpha_{i,k}^{(t)}}+\alpha_{j,l}^{2}\dfrac{\alpha_{i,k}^{(t)}}{2\alpha_{j,l}^{(t)}}\right)\log p_{k,l}(y_{i,j})\vspace{0.25cm}\\ &+&\displaystyle\sum\limits_{i=1}^{N}\displaystyle\sum\limits_{k=1}^{K}\alpha_{i,k}\left(\log\gamma_{k}^{(t)}-\log\alpha_{i,k}^{(t)}-\dfrac{\alpha_{i,k}}{\alpha_{i,k}^{(t)}}+1\right)\vspace{0.25cm}\\ &=&\displaystyle\sum\limits_{i=1}^{N}\displaystyle\sum\limits_{k=1}^{K}\dfrac{\Omega_{i,k}^{(t)}(\mathbf{y},\bm{\alpha}^{(t)},\bm{\theta}^{(t)})}{\alpha_{i,k}^{(t)}}\alpha_{i,k}^{(t)}+\displaystyle\sum\limits_{i=1}^{N}\displaystyle\sum\limits_{k=1}^{K}\alpha_{i,k}\left(\log\gamma_{k}^{(t)}-\log\alpha_{i,k}^{(t)}-\dfrac{\alpha_{i,k}}{\alpha_{i,k}^{(t)}}+1\right)\vspace{0.25cm}\\ &=&\displaystyle\sum\limits_{i=1}^{N}\displaystyle\sum\limits_{k=1}^{K}\left(\dfrac{\Omega_{i,k}^{(t)}(\mathbf{y},\bm{\alpha}^{(t)},\bm{\theta}^{(t)})}{\alpha_{i,k}^{(t)}}-\dfrac{1}{\alpha_{i,k}^{(t)}}\right)\alpha_{i,k}^{2}+\left(\log\gamma_{k}^{(t)}-\log\alpha_{i,k}^{(t)}+1\right)\alpha_{i,k}\vspace{0.25cm}\\ &=&\displaystyle\sum\limits_{i=1}^{N}\displaystyle\sum\limits_{k=1}^{K}A_{i,k}(\mathbf{y},\bm{\alpha}^{(t)})\alpha_{i,k}^{2}+B_{i,k}(\mbox{\boldmath$\gamma$}^{(t)},\bm{\alpha}^{(t)})\alpha_{i,k}\end{array}$

### v2

@@ 4.1 Signed Stochastic Block Model Approximation
For any model specification, we recover the Signed Stochastic Block Model (SSBM), as shown in Example 1, by setting all parameters in model (4) corresponding to dyad-dependent statistics (e.g., transitivity) to zero. Note that the sufficient statistics in (3) are already dyad-independent by definition. Therefore, the approximation of the probability of our model by the probability of an SSBM only accrues errors for within-block networks. As $N$ and $K$ grow, our joint probability will then increasingly be dominated by terms relating to between-block probabilities, provided single blocks are not too large. This offers a heuristic justification for why the SSBM provides a good approximation in large populations.
@@
All sufficient statistics considered in this paper satisfy Assumptions 4.1 and 4.1. Following Theorem 2 from Babkin et al. (2020), the full conditional probability $\mathbb{P}_{\bm{\theta}}(\mathbf{Y}=\mathbf{y}\mid\bm{Z}=\bm{z})$ can then be approximated with $\bm{\theta}^{\text{SBM}}$ as follows
@@ 4.2 Variational Approximation based on MM Updates
Supported by this result, we substitute $\mathbb{P}_{\bm{\theta}}(\mathbf{Y}=\mathbf{y}\mid\bm{Z}=\bm{z})$ in (8) by the probability of an SSBM, which simplifies the optimization in Steps 1 and 2 outlined above. We assume a mean-field approximation of $\mathbb{P}(\bm{Z}=\bm{z}\mid\mathbf{Y}=\mathbf{y})$, implying that the joint distribution of $\bm{Z}$ factorizes across nodes:
@@
4.3 Estimation with known Blocks
@@
$\begin{array}[]{llllllllll}\hat{\bm{z}}=\left(\operatorname*{arg\,max}\limits_{k\,\in\,\{1,\ldots,K\}}\alpha_{i,k}\right)_{i=1}^{N}.\end{array}$
@@
The term $\bm{a}\otimes\bm{b}$ denotes the Kronecker product of vectors $\bm{a}\in\mathbb{R}^{m}$ and $\bm{b}\in\mathbb{R}^{n}$. In the following, we focus on estimating the within-block model parameter $\bm{\beta}_{w,\operatorname{vec}}$. The method applies analogously to $\bm{\beta}_{b,\operatorname{vec}}$. While both MCMC-based maximum likelihood (Hunter and Handcock, 2006) and maximum pseudo-likelihood methods (MPLE, Strauss and Ikeda, 1990) can be used to estimate the parameters in the second step, MPLE is generally better suited for large networks due to its computational efficiency and scalability. If dyadic independence holds, both approaches coincide.
@@
With $\Delta_{i,j,k}^{0\rightarrow+}(\mathbf{y})$ and $\Delta_{i,j,k}^{0\rightarrow-}(\mathbf{y})$ being the change in the reparametrized sufficient statistics given in (11) after changing the edge value $y_{i,j}$ from ``$0"$ to ``+"``"``+"` ` + " or ``"``"``-"` ` - ", respectively, we can express the relative log-odds of observing $\mathbf{Y}_{i,j}$ to be ``+"``"``+"` ` + " abs ``"``"``-"` ` - " instead of ``$0"$, conditional on $z_{i}=z_{j}=k$:
$\begin{array}[]{llllllllll}\log\left(\dfrac{\mathbb{P}_{\bm{\beta}_{w,\operatorname{vec}}}(Y_{i,j}=``+"\mid\mathbf{Y}_{(-ij)}=\mathbf{y}_{(-ij)},z_{i}=z_{j}=k)}{\mathbb{P}_{\bm{\beta}_{w,\operatorname{vec}}}(Y_{i,j}=``0"\mid\mathbf{Y}_{(-ij)}=\mathbf{y}_{(-ij)},z_{i}=z_{j}=k)}\right)&=&\bm{\beta}_{w,\operatorname{vec}}^{\top}\Delta_{i,j,k}^{0\rightarrow+}(\mathbf{y})\vspace{0.25cm}\\ \log\left(\dfrac{\mathbb{P}_{\bm{\beta}_{w,\operatorname{vec}}}(Y_{i,j}=``-"\mid\mathbf{Y}_{(-ij)}=\mathbf{y}_{(-ij)},z_{i}=z_{j}=k)}{\mathbb{P}_{\bm{\beta}_{w,\operatorname{vec}}}(Y_{i,j}=``0"\mid\mathbf{Y}_{(-ij)}=\mathbf{y}_{(-ij)},z_{i}=z_{j}=k)}\right)&=&\bm{\beta}_{w,\operatorname{vec}}^{\top}\Delta_{i,j,k}^{0\rightarrow-}(\mathbf{y}).\end{array}$
@@
Newton-Raphson algorithms are commonly used to maximize (12) based on (13) with the following gradient and negative Hessian:
@@ 4.4 Uncertainty Quantification
Given the previously sampled block memberships, $\bm{z}^{(1)},\ldots,\bm{z}^{(T)}$, the parameters of the model,$(\bm{\beta}_{w,\operatorname{vec}}^{(1)},\mbox{\boldmath$\Sigma$}^{(1)}),\ldots,(\bm{\beta}_{w,\operatorname{vec}}^{(T)},\mbox{\boldmath$\Sigma$}^{(T)})$, are re-estimated for each sampled partition to obtain valid uncertainty quantification. The estimated parameters $\bm{\beta}_{w,\operatorname{vec}}^{(t)}$ are pooled as follows
@@ 5 Simulation Study
This section evaluates the performance of the proposed estimation procedure in both steps. For the evaluation of the parameter recovery in Step 2, we use MPLE as they are more scalable than MCMC-based methods. To evaluate the block recovery, we use the variational approach described in Section 4 and compare it to the spectral clustering described in Lei and Rinaldo (2015), for which we treat the network as a binary one. The block recovery will be assessed using the Yule’s $\phi$-coefficient:
@@ 6 Wikipedia Network
We demonstrate the proposed model in an application to a network of Wikipedia editors who contribute to and refine content across a range of topics. Wikipedia is a free, community-edited online encyclopedia that has been the subject of numerous prior studies (Brandes et al., 2009; Iba et al., 2010). A subset of these editors can be considered experts in specific fields, as they focus their contributions to a limited number of Wikipedia pages within their area of expertise. These editors typically exhibit limited awareness of edits or interactions occurring outside their domain. Such behavior gives rise to block structure and localized dependence within the network, where each block corresponds to a distinct area of knowledge.
@@ 6.1 Model Specification
Figure 3: Out-of-sample cross-validation results for the “Full Triad” model including all triadic terms ($\text{Edges}^{+/-}$, $\text{GWD}^{+/-}$, and all GWESP). The distribution of simulated statistics across 100 replications is compared against the observed statistics for each block.
@@
$\begin{array}[]{llllllllll}\bm{s}(\mathbf{y}_{k,k})&=&(&\text{Edges}^{+}(\mathbf{y}_{k,k}),\,\text{Edges}^{-}(\mathbf{y}_{k,k}),\,\text{GWD}^{+}(\mathbf{y}_{k,k},0.2),\,\text{GWD}^{-}(\mathbf{y}_{k,k},0.2),\\ &{}&{}&\text{GWESE}^{+}(\mathbf{y}_{k,k},0),\,\text{GWESF}^{+}(\mathbf{y}_{k,k},0),\,\text{GWESE}^{-}(\mathbf{y}_{k,k},0),\\ &{}&{}&\text{GWESF}^{-}(\mathbf{y}_{k,k},0))^{\top}.\end{array}$
The results of the out-of-sample cross-validation, shown in Figure 3 and Supplement C.1, suggest that the “Full Triad” model, including edges, GWD and all GWESP terms, is better equipped to describe the Wikipedia network.
@@ 6.2 Results
The edges terms are modeled as a function of block size $\log(N_{k})$. The terms $\text{Edges}^{+}$ and $\text{Edges}^{-}$ represent the intercepts, which alone are not directly interpretable, as they would only apply to hypothetical blocks of size 1. Instead, the combination of intercept and block size effect reflects the expected edge probability within a given block. Across all models, the coefficients for positive edges are lower than for negative edges, indicating that negative edges are more frequent within blocks. The consistently negative coefficients for the block size terms demonstrates that larger blocks tend to be sparser, with edge probability decreasing as block size increases. This reflects the fact that the number of possible edges grows quadratically, while the number of actual edges grows at a slower pace. Adding GWD terms reduces the edge coefficients, suggesting that some variation in edge probability is explained by degree structure. The GWD coefficients are negative for both edge types in all models, indicating a general tendency against high-degree nodes. The effect is stronger for positive edges, implying that these edges are more evenly distributed and less likely to form hubs. As expected from structural balance theory, the $\text{GWESE}^{+}$ terms have positive and statistically significant coefficients in both the “Full Triad” and “Partial Triad” model. This means that users are more likely to restore the work of another user if they share a common enemy. While still positive and significant, the other balanced triad represented by $\text{GWESF}^{+}$ is closer to zero, similar to the unbalanced triad $\text{GWESE}^{-}$. However, the term $\text{GWESF}^{-}$ is significantly larger than $\text{GWESF}^{+}$, which is not consistent with structural balance theory. This discrepancy suggests that structural balance theory may not fully apply to this expert network, For instance, undoing a friend-of-a-friend’s contribution could be more common due to task-related disagreements rather than group-based antagonism. Overall, the more complex models capture important network features, such as triadic relationships, and this is reflected in their lower AIC values, indicating better model fit despite increased complexity.
Our Wikipedia network analysis of thousands of subject-matter experts serves an example where global dependence assumption seems unrealistic. Through the use of local dependence, we found some adherence to the behavioral norms predicted by structural balance theory. Specifically, users who have an enemy in common are more likely to restore each other’s contributions. This finding is consistent with Lerner and Lomi (2019).
@@ 7 Conclusion
Our proposed model for signed networks combines the strengths of stochastic block models and exponential random graph models, while being scalable to thousands of nodes. There are several directions for extending the proposed model. First, we assumed that the number of blocks, $K$, was known a priori. In practice, this may not hold, and methods to infer $K$ from data would enhance model flexibility (Saldaña et al., 2017). Second, the assumption of nonoverlapping community structure could be relaxed by allowing nodes to belong to multiple blocks, as in mixed-membership models (see, e.g., Latouche et al., 2011). This would accommodate more realistic block structures. Third, the observed degree heterogeneity in the Wikipedia network highlights the need for a degree-corrected stochastic block model for signed networks.
@@ B.1 Surrogate Function
$\begin{array}[]{llllllllll}Q({\mbox{\boldmath$\gamma$}}^{(t)},{\bm{\theta}}^{(t)};{\bm{\alpha}}^{(t)},{\bm{\alpha}})&=&\displaystyle\sum\limits_{i<j}\displaystyle\sum\limits_{k=1}^{K}\displaystyle\sum\limits_{l=1}^{K}\left(\alpha_{i,k}^{2}\dfrac{\alpha_{j,l}^{(t)}}{2\alpha_{i,k}^{(t)}}+\alpha_{j,l}^{2}\dfrac{\alpha_{i,k}^{(t)}}{2\alpha_{j,l}^{(t)}}\right)\log p_{k,l}(y_{i,j})\vspace{0.25cm}\\ &+&\displaystyle\sum\limits_{i=1}^{N}\displaystyle\sum\limits_{k=1}^{K}\alpha_{i,k}\left(\log\gamma_{k}^{(t)}-\log\alpha_{i,k}^{(t)}-\dfrac{\alpha_{i,k}}{\alpha_{i,k}^{(t)}}+1\right)\vspace{0.25cm}\\ &=&\displaystyle\sum\limits_{i=1}^{N}\displaystyle\sum\limits_{k=1}^{K}\dfrac{\Omega_{i,k}^{(t)}(\mathbf{y},\bm{\alpha}^{(t)},\bm{\theta}^{(t)})}{\alpha_{i,k}^{(t)}}\alpha_{i,k}^{(t)}\vspace{0.25cm}\\ &+&\displaystyle\sum\limits_{i=1}^{N}\displaystyle\sum\limits_{k=1}^{K}\alpha_{i,k}\left(\log\gamma_{k}^{(t)}-\log\alpha_{i,k}^{(t)}-\dfrac{\alpha_{i,k}}{\alpha_{i,k}^{(t)}}+1\right)\vspace{0.25cm}\\ &=&\displaystyle\sum\limits_{i=1}^{N}\displaystyle\sum\limits_{k=1}^{K}\left(\dfrac{\Omega_{i,k}^{(t)}(\mathbf{y},\bm{\alpha}^{(t)},\bm{\theta}^{(t)})}{\alpha_{i,k}^{(t)}}-\dfrac{1}{\alpha_{i,k}^{(t)}}\right)\alpha_{i,k}^{2}\vspace{0.25cm}\\ &+&\displaystyle\sum\limits_{i=1}^{N}\displaystyle\sum\limits_{k=1}^{K}\left(\log\gamma_{k}^{(t)}-\log\alpha_{i,k}^{(t)}+1\right)\alpha_{i,k}\vspace{0.25cm}\\ &=&\displaystyle\sum\limits_{i=1}^{N}\displaystyle\sum\limits_{k=1}^{K}A_{i,k}(\mathbf{y},\bm{\alpha}^{(t)})\alpha_{i,k}^{2}+B_{i,k}(\mbox{\boldmath$\gamma$}^{(t)},\bm{\alpha}^{(t)})\alpha_{i,k}\end{array}$